OpenStudy (anonymous):

What is the value of 12C5 ? A. 95,040 B. 792 C. 120 D. 60

OpenStudy (anonymous):

I got B. I plugged the values into the formula: \[C(n,r)=\frac{ n! }{(n-r)!r! }\]

OpenStudy (anonymous):

so that would be:\[C(12,5)=\frac{ 12! }{ 5!7! }\]

OpenStudy (anonymous):

ohhhh.... ok thanks i might have a couple more questions

OpenStudy (anonymous):

What is the value of (8 – 3)!? A. 5 B. 120 C. 40,314 D. 336 My answer is a am i right?

OpenStudy (anonymous):

hold on i'll figure it out...

OpenStudy (anonymous):

It's B because (8-3)! means the factorial of 5. To figure it out you'd multiply 5*4*3*2*1.

OpenStudy (anonymous):

okay thanks, umm could you answer one more

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Troy can order a large or small snow cone. There are 15 flavors and 2 toppings from which to choose. How many ways can Troy order a snow cone? A. 30 B. 60 C. 32 D. 120 I'm thinking b

OpenStudy (anonymous):

Yeah, I think that's it because 2 flavors*2 sizes*15 flavors=60 options

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

you're welcome. any more questions before I go?

OpenStudy (anonymous):

lol i was hoping you would ask, actually i do have 1 or 2 idk

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

What is the value of 7P3? A. 35 B. 210 C. 56 D. 840

OpenStudy (anonymous):

So for this one we'd use the formula:\[P(n,r)=\frac{ n! }{ (n-r)!}\] So we'd plug the values in and get 210.

OpenStudy (anonymous):

okayyyyyy I just need you to check this one A book club has 10 members. How many ways can they choose a president, vice president, and secretary from among the members? A. 30 B. 120 C. 720 D. 100 i pick c

OpenStudy (anonymous):

Yeah, I think that's it because they'd have 10 choices for president, then 9 choices for vice president, then 8 choices for secretary, so if you'd multiply those you would get 720.

OpenStudy (anonymous):

ok now i have a question that has nothing to do with learning

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

are you a girl or boy?

OpenStudy (anonymous):

I'm a girl.

OpenStudy (anonymous):

oh okay, I mastered my quiz so thanks! :)

OpenStudy (anonymous):

you're welcome!